OpenStudy (anonymous):

solve for x x - 7/8= 7 /2

OpenStudy (anonymous):

@hartnn

OpenStudy (diyadiya):

is it \[\frac{ x-7 }{ 8 }=\frac{7}{2}\] or \[x-\frac{ 7 }{ 8 }=\frac{ 7 }{2 }\]

OpenStudy (anonymous):

the top one.

OpenStudy (diyadiya):

1.multiply both sides by 8 2. multiply both sides by 2

OpenStudy (anonymous):

how would i do that?

OpenStudy (diyadiya):

\[\frac{x-7}{8} \times 8 = \frac{7}{2} \times 8\] 8 will get cancelled \[x-7 = \frac{56}{2}\] Now multiply both sides by 2\[(x-7)\times2 = \frac{56}{2} \times 2\] \[2(x-7) =56\]

OpenStudy (anonymous):

ok whats next?

OpenStudy (diyadiya):

\[2x-14=56\]

OpenStudy (diyadiya):

Now add 14 on both sides

OpenStudy (diyadiya):

what will you get?

OpenStudy (anonymous):

where should i add 14 to? it may seem stupid but im having a hard time with this.

OpenStudy (anonymous):

70! right

OpenStudy (anonymous):

?

OpenStudy (diyadiya):